OpenStudy (anonymous):

all vectors of the form (a,b,c) where b=a+c is this subspace of R^3

OpenStudy (anonymous):

A subspace S must satisfy the three conditions: \[0\in S\] \[\lambda u \in S\] \[u +v \in S\] when u en v are in S.

OpenStudy (anonymous):

It is A subspace but it is obviously not "the" subspace.

OpenStudy (anonymous):

Let's call your subspace S. (0,0,0) satisfies 0=0+0, so the first condition is true. when (a,b,c) is in S then b=a+c, then is also true: \[\lambda b=\lambda a+\lambda\] So the second condition is satisfied

OpenStudy (anonymous):

\[b _{u}=a _{u}+c _{u}\] \[b _{v}=a _{v}+c _{v}\] \[b _{v}+b _{u}=a _{v}+a _{u}+c _{v}+c _{u}\] Which satisfies the third condition, so it is indeed a subspace of R^3

OpenStudy (anonymous):

v = (a_1 + a_2, b_1 + b_2, c_1 + c_2) b_1 = a_1 + c_1 b_2 = a_2 + c_2 b_1 + b_2 = a_1 + c_1 + a_2 + c_2 (b_1 + b_2) = (a_1 + a_2) + (c_1 + c_2)

OpenStudy (anonymous):

Also it should be mentioned that the three mentioned properties are: Existence of a zero vector, Closure under addition Closure under multiplication

OpenStudy (anonymous):

i got the concept thanks sir